OpenStudy (anonymous):

the sum of n terms of two arithmetic progression are in the ratio (3n+8):(7n+15). find the ratio of their 12th terms?

OpenStudy (anonymous):

infinite

OpenStudy (compassionate):

WELCOME TO OPEN STUDY!

OpenStudy (anonymous):

(2001+1)x1000 + 1001

OpenStudy (anonymous):

a = 1 l = 2001 n = 2001 use the arithmetic series formula

OpenStudy (anonymous):

Come on, everyone knows this? 1 + 2 +...............+2000 +2001 2001 + 2000 +...........+2 + 1

OpenStudy (compassionate):

There are 1000 pairs of numbers that add up to 2001 1000*2001 = 2,001,000 In general, the sum of the consecutive integers from a to b = (a+b)*(b-a+1)/2

OpenStudy (lgbasallote):

use the formula for sum of arithmetic sequence \[S_n = \frac{n(a_1 + a_{n})}{2}\] where Sn is the sum n is the number of terms a1 is the first term an is the last term since the last term is 2001 we can say there are 2001 terms so let's substitute \[S_n = \frac{2001(1 + 2001)}{2}\] \[S_n = \frac{2001(2002)}{2}\] \[S_n = 2001 \times 1001\] use a calculator now i cannot solve this mentally

OpenStudy (anonymous):

1 + 2 + 3 .......................n The sum to n terms is given by: \(Sum = \frac{n(n+1)}{2}\) Use this to find Sum.. n = 2001 here..

OpenStudy (compassionate):

ANS: 2,003,001

OpenStudy (anonymous):

thanks

OpenStudy (compassionate):

Best of luck, Bas.

OpenStudy (anonymous):

\[\large \color{green}{\textbf{Welcome To Openstudy..}}\]

OpenStudy (anonymous):

Do not edit the question just click on close button and then post the new question.. @basiljohn